Write a formula for the function g(x) that results when the graph of f(x)=x2     is horizontally stretched by a factor of 3, then shifted to the left 4 units, and then shifted down 3 units . Sketch the graph of f(x)  fas well as the graph of the new transformed function g(x)   

g(x)=? 



1
Expert's answer
2021-05-18T08:12:07-0400

"f(x)=x^2.\newline
\text{horizontally stretched by a factor of 3.}\newline
a(x)=f(\frac{x}{3})=\frac{x^2}{9}\newline\text{
shifted to the left 4 units,}\newline
b(x)=a(x+4)=f(\frac{x+4}{3})=\frac{(x+4)^2}{9}\newline\text{
shifted down 3 units,}\newline
g(x)=b(x)-3=a(x+4)-3=f(\frac{x+4}{3})-3=\frac{(x+4)^2}{9}-3\newline\text{
Then, the required }g(x)=\frac{(x+4)^2}{9}-3.\newline"